Calculate the real wage(W/P) if the value of the markup, u=0.1


1
Expert's answer
2022-02-01T11:32:28-0500

Calculate the real wage(W/P) if the value of the markup, u=0.1


Price setting equals the wage-setting equation via the real wage:

"\frac{w}{p}=\frac{1}{1+m}=1-u"

Where m is markup and u is the unemployment rate.

Real wage = "\frac{w}{p}=\frac{1}{1+m}=\frac{1}{0.1}"

= "\frac{1}{1.1}"

Therefore, real wage "\frac{w}{p}" = 0.9091